OpenStudy (anonymous):

Solve each of the following systems of equations by substitution. a. 2x + 10y = 36 y=x + 12 b. 3x + 5y = 33 x = 2y c. 10x - 10y = 65 x = 10y + 2 d. 5x + y = 18 x= y

OpenStudy (amistre64):

2x + 10y = 36 [ y=x + 12] -10 2x + 10y = 36 -10y=-10x - 120 --------------------- 2x = -10x -84 12x = -84 x = -84/12 = -42/6 = -7

OpenStudy (amistre64):

y = (-7) + 12; y = 5

OpenStudy (anonymous):

okay